# BOLIMES: Boruta–LIME optiMized fEature Selection for Gene Expression Classification

### v1

@@ 1 Introduction
To address these challenges, feature selection plays a critical role in gene expression classification. By identifying and retaining only the most informative genes, feature selection not only improves model generalization but also enhances interpretability, allowing researchers to derive biologically meaningful insights from machine learning predictions [9, 20]. Furthermore, reducing the feature space significantly lowers computational costs, enabling faster model training and inference. More importantly, gene selection aligns with the biological reality that only a subset of genes actively contributes to disease mechanisms, making feature selection a crucial step in biomedical analysis. Without an effective selection strategy, classifiers are prone to noise, reduced accuracy, and difficulty in extracting meaningful biomarkers for clinical applications.
@@
To bridge this gap, we propose BOLIMES (Boruta–LIME Enhanced eXtraction), a novel feature selection algorithm designed to enhance gene expression classification by systematically refining the feature subset. Unlike conventional methods that rely solely on statistical ranking or classifier-specific selection, BOLIMES integrates the robustness of Boruta [24, 36] with the interpretability of LIME [30], ensuring that only the most relevant and influential genes are retained. BOLIMES first applies Boruta, a powerful wrapper method that iteratively refines feature sets by comparing real genes against randomly permuted shadow features, ensuring that no informative gene is prematurely discarded. Then, LIME [30, 23] (Local Interpretable Model-agnostic Explanations) is employed to assess the local importance of each selected gene, providing an interpretable ranking of features that capture model-specific contributions. Finally, an iterative classification evaluation determines the optimal feature subset by selecting the number of genes that maximize predictive accuracy. By integrating Boruta’s exhaustive feature selection with LIME’s interpretability-driven refinement, BOLIMES effectively balances dimensionality reduction with model performance, making it a powerful solution for high-dimensional gene expression analysis.
The remainder of this paper is structured as follows: we briefly presents a fundamental background in Section 2. Next, Section 3 describes our proposal algorithm. Then, Section 4 represents the experiment, the results of the models and an discussion. Finally, Section 5 shows the conclusions and future work.
@@ 2 Background
In this section, we present the foundation of our approach with integrating Boruta and Lime. We also provide the ML algorithm that will be used for the gene expression classification task.
@@ 2.1 Gene Expression Classification with ML models
While numerous feature selection techniques exist, our study concentrates on two: Boruta and LIME. We choose Boruta for its robust ability to identify all truly relevant features within high-dimensional datasets, ensuring that no significant predictor is overlooked. Complementing this, LIME is employed for its exceptional capacity to provide local, interpretable explanations of model predictions, which is crucial when assessing feature importance through numerous distance calculations. Together, these methods offer a powerful and efficient approach to refining our gene expression classification models. Now, we introduce Boruta and Lime in the next session.
@@ 2.2 Leveraging Boruta for Robust Feature Extraction
Boruta [24, 36] is a powerful wrapper-based feature selection algorithm designed to identify all truly relevant variables in a dataset. By comparing the importance of actual features with that of randomly generated “shadow” features, Boruta systematically filters out irrelevant variables while preserving essential predictors. This rigorous selection process is particularly valuable in high-dimensional applications, such as gene expression classification, where capturing meaningful signals is crucial. Rather than directly improving predictive accuracy, Boruta refines the feature set, which can indirectly enhance model interpretability and performance. For clarity, we formally define Boruta as follows:
@@
    Iteration: Remove irrelevant features and repeat until all features are decisively classified.
@@ 2.3 XAI for Feature Selection
Explainable AI (XAI) [12, 35] represents a forefront of AI research, aiming to elucidate the decision-making processes of complex models. In the context of gene expression classification, where feature selection is pivotal to model performance and interpretability, our study leverages LIME—Local Interpretable Model-Agnostic Explanations—to demystify and extract critical features. LIME approximates the behavior of a sophisticated, black-box model with a simpler, locally interpretable surrogate, thereby pinpointing the most influential predictors in the vicinity of a given instance. This approach enhances the transparency of the model’s predictions andfacilitates a more informed and rigorous feature selection process, ultimately contributing to both improved accuracy and trustworthiness of the classification system. Now, we provide a formal definition of LIME as follows:
@@
3 BOLIMEX algorithm
Our methodology begins with applying the Boruta algorithm to sift through the high-dimensional gene expression dataset, effectively filtering out irrelevant features and isolating those that are truly significant. This initial reduction is critical because LIME, our subsequent interpretability tool, involves generating numerous perturbed samples and calculating distances—a computationally intensive process, especially in large feature spaces. By narrowing the focus to a refined subset via Boruta, we significantly reduce the computational burden and enhance the precision of LIME’s local explanations. This sequential approach streamlines the overall feature selection process and fortifies the reliability and clarity of our model’s interpretability, ultimately leading to improved classification performance.
@@ 14                   $f_{\text{opt}}\leftarrow f_{k}$
The algorithm presents an optimal feature selection framework by integrating Boruta and LIME to refine feature subsets for classification. Given a dataset $D=(X,y)$, it first applies Boruta to extract a subset $X^{*}$ of relevant features. LIME then computes local importance scores $\mathcal{I}$ for $X^{*}$, producing a ranked set $X^{*}_{R}$ where
@@
The best-performing classifier defines $k^{*}$, yielding the final subset $X_{\text{opt}}$ and trained model $f_{\text{opt}}$.
@@ 4 Experiment and results
This section offers a brief description of the gene expression datasets while delivering a detailed comparative analysis of the classification models. Furthermore, we also provide the results of Boruta. Our source code has been made publicly accessible on GitHub [^1] https://github.com/Vannguyen0312/Manholes-App/.
@@ 4.1 Dataset and Configurations
The gene expression datasets summarized in Table 1 exemplify the inherent challenges of high-dimensional biomedical data. With sample sizes ranging from 53 to 575 and feature counts spanning from approximately 11,950 to over 54,600, these datasets present a significant imbalance between the number of available samples and the vast dimensionality of gene expression profiles. Additionally, the variability in the number of classes—from as few as 3 to as many as 14—further complicates the classification task by introducing diverse and complex biological signals. This high dimensionality coupled with limited sample sizes accentuates the risk of overfitting and underscores the critical need for effective feature selection. Robust feature selection methods are essential to isolate the most informative genes, thereby enhancing model interpretability and predictive accuracy in gene expression classification. In this study, our feature selection strategy is specifically designed to address these challenges, ensuring that only the most relevant features are retained for subsequent classification tasks.
@@ 15 E-GEOD-44077 226 33252 4
To evaluate our approach and the training model, we implement the AI libraries (i.e., tensorflow (version 3.10.0), Keras (version 3.0)) and run the experiments on the computer with the following configuration: AMD Ryzen 7 6800HS, Ceator Edition 3.20 GHZ, 16 GB RAM, Windows 10 OS.
@@ 4.2 Boruta’s Feature Selection
In this study, we adopt Boruta as our primary feature selection method to substantially reduce the dimensionality of our gene expression datasets before applying LIME for model interpretation. The rationale behind this choice is to prevent an explosion in computational complexity and potential loss of interpretability when LIME is applied to an excessively high-dimensional feature space. By using Boruta, we are able to effectively eliminate irrelevant features while retaining those that are truly informative for classification. As evidenced in Table 2, datasets such as E-GEOD-20685 were reduced from 54,627 dimensions to only 545 confirmed features, thereby rendering the subsequent LIME analysis both feasible and efficient. To achieve this, Boruta was executed with the following parameters via the BorutaPy library: a random forest classifier (‘rf‘) with 300 estimators, a maximum of 200 iterations, an alpha value of ’0.01’, a percentile threshold of 95, two-step feature selection enabled, and a fixed random state of 42, with verbose output enabled. These parameter settings were meticulously chosen to ensure a rigorous and robust selection process, ultimately facilitating a more interpretable and high-performing classification model.
@@ 4.3 Classification results
In our study, we conducted extensive experiments with four advanced algorithms to determine the method yielding the highest accuracy across various gene expression datasets. We configured an SVM using SVC with a radial basis function kernel, setting C to 100,000, gamma to ”scale”, and class weights to ”balanced”, ensuring reproducibility with a random state of 42. In parallel, we employed a Random Forest classifier with 200 estimators, a maximum depth of 10, and balanced class weights. Additionally, we implemented both an XGBClassifier and a GradientBoostingClassifier, each configured with 50 estimators, a maximum depth of 10, and a learning rate of 0.01, with the random state fixed at 42. Our comparative evaluation revealed that while the SVM excelled on datasets with fewer samples, the ensemble methods—particularly Random Forest and XGBoost—demonstrated more robust performance on high-dimensional data. This comprehensive analysis highlights the critical role of algorithm selection and fine-tuning in achieving optimal classification performance in gene expression studies.
@@
ID Dataset Class Method Samples Top K Classification Results Time (s)
Features Acc Prec Rec F1 Score Training Preprocess Total
ours 327 107 0,955 0,956 0,955 0,955 1,151 12,951 112,137
1 20685 6 327 107 0,955 0,956 0,955 0,955 1,151 12,951 112,137
327 107 0,955 0,956 0,955
2 20711 5 90 67 0,944 0,952 0,944 0,944 0,298 3,745 17,754
3 21050 4 310 70 0,677 0,710 0,677 0,678 1,494 10,400 295,982
4 21122 7 158 170 0,938 0,915 0,938 0,922 0,548 2,010 82,092
5 29354 3 53 10 1,000 1,000 1,000 1,000 0,165 0,823 0,165
6 30784 3 229 14 1,000 1,000 1,000 1,000 0,730 8,197 3,476
7 31312 3 498 43 0,900 0,900 0,900 0,900 0,063 19,236 2,144
8 31552 3 111 15 0,957 0,917 0,957 0,936 0,299 2,642 1,816
9 32537 7 217 40 0,796 0,730 0,796 0,747 0,735 0,918 21,201
10 33315 10 575 273 0,878 0,860 0,878 0,867 0,138 4,844 33,250
11 36895 14 76 15 0,929 0,893 0,929 0,905 0,228 3,199 1,304
12 37364 4 94 17 0,947 0,965 0,947 0,947 0,259 3,393 2,083
13 39582 6 566 251 0,877 0,889 0,877 0,880 0,130 20,512 25,058
14 39716 3 53 12 1,000 1,000 1,000 1,000 0,157 1,642 0,454
15 44077 4 226 18 1,000 1,000 1,000 1,000 0,578 5,513 5,220
The classification results obtained using SVM, Random Forest, XGBoost, and Gradient Boosting reveal distinct performance and computational trade-offs. SVM generally delivers high accuracy and balanced precision, recall, and F1 scores across most datasets, with relatively low training times in some cases; however, its preprocessing time can be substantial when handling a large number of features. In contrast, Random Forest exhibits competitive accuracy, often outperforming SVM on certain datasets, yet it tends to incur significantly higher training times, particularly when the number of top features increases—as observed in dataset 1 where training time was notably high. XGBoost demonstrates robust performance with competitive accuracy and precision; its moderate training and preprocessing times suggest that it strikes a balance between model complexity and computational efficiency, though its performance can be dataset-dependent (e.g., lower accuracy on dataset 13). Meanwhile, Gradient Boosting achieves near-perfect accuracy on several datasets, indicating high robustness in classification, yet its training and total processing times vary considerably across datasets.
@@
ID Dataset Class Samples Top N Classification Results Time (s)
Features Acc Prec Rec F1 Score Training Preprocess Total
1 20685 6 327 487 0,970 0,971 0,970 0,970 40,762 13,055 16306,229
2 20711 5 90 10 0,833 0,841 0,833 0,830 5,396 3,703 5,396
3 21050 4 310 72 0,726 0,727 0,726 0,711 25,093 10,252 1507,624
4 21122 7 158 104 0,938 0,922 0,938 0,925 11,577 1,971 1039,209
5 29354 3 53 10 1,000 1,000 1,000 1,000 3,325 1,059 3,325
6 30784 3 229 12 1,000 1,000 1,000 1,000 14,531 8,278 43,917
7 31312 3 498 161 0,900 0,821 0,900 0,858 1,594 19,739 221,322
8 31552 3 111 12 1,000 1,000 1,000 1,000 6,667 2,608 20,318
9 32537 7 217 87 0,773 0,680 0,773 0,707 17,146 0,935 1238,228
10 33315 10 575 444 0,887 0,854 0,887 0,868 3,215 4,947 985,866
11 36895 14 76 20 0,929 0,929 0,929 0,929 4,347 3,242 46,355
12 37364 4 94 10 0,947 0,953 0,947 0,947 5,579 3,380 5,579
13 39582 6 566 189 0,877 0,890 0,877 0,876 1,908 20,970 304,137
14 39716 3 53 15 1,000 1,000 1,000 1,000 3,172 1,528 19,338
15 44077 4 226 52 1,000 1,000 1,000 1,000 15,408 6,125 642,643

### v2

@@ 1 Introduction
To address these challenges, feature selection [28, 30] plays a critical role in gene expression classification. By identifying and retaining only the most informative genes, feature selection not only improves model generalization but also enhances interpretability, allowing researchers to derive biologically meaningful insights from machine learning predictions [10, 21]. Furthermore, reducing the feature space significantly lowers computational costs, enabling faster model training and inference. More importantly, gene selection aligns with the biological reality that only a subset of genes actively contributes to disease mechanisms, making feature selection a crucial step in biomedical analysis. Without an effective selection strategy, classifiers are prone to noise, reduced accuracy, and difficulty in extracting meaningful biomarkers for clinical applications.
@@
To bridge this gap, we propose BOLIMES (Boruta–LIME Enhanced eXtraction), a novel feature selection algorithm designed to enhance gene expression classification by systematically refining the feature subset. Unlike conventional methods that rely solely on statistical ranking or classifier-specific selection, BOLIMES integrates the robustness of Boruta [25, 39] with the interpretability of LIME [33], ensuring that only the most relevant and influential genes are retained. BOLIMES begins by applying Boruta, safeguarding against the premature elimination of any informative genes. Next, LIME (Local Interpretable Model-agnostic Explanations) [33, 24] is employed to assess the local importance of each selected gene, offering an interpretable ranking that highlights the contributions of individual features to the model. Finally, an iterative classification evaluation identifies the optimal subset of features by selecting the number of genes that maximize predictive accuracy. We pay attention at Boruta and LIME for the following reasons: (1) Due to the large number of dimensions, LIME would face scalability issues when it comes to interpretation. (2) Boruta is viewed as a global feature selection method, whereas LIME focuses on local feature selection. Therefore, we will first perform global filtering using Boruta, followed by local refinement with LIME. (3) Our primary objective is to determine the optimal number of features for selection. As a result, we have proposed the Bolimex algorithm to select the most appropriate set of features, ensuring the best possible accuracy.
The remainder of this paper [^1] This paper will be condensed to meet the page limit requirements of the conference upon acceptance. We aim to retain all the core ideas and details from this version to ensure the integrity of our research. is structured as follows: we briefly presents a fundamental background in Section 2. Next, Section 3 describes our proposal algorithm. Then, Section 4 represents the experiment, the results of the models and an discussion. Finally, Section 5 shows the conclusions and future work.
@@ 2 Background
In this section, we present the foundation of our approach with integrating Boruta and Lime. We also provide the ML models that will be used for the gene expression classification task.
@@ 2.1 Gene Expression Classification with ML models
Feature selection is crucial before classification begins. Our study focuses on two techniques: Boruta and LIME. We now introduce Boruta and LIME in the following sections.
@@ 2.2 Leveraging Boruta for Robust Feature Extraction
Boruta [25, 39] is a powerful wrapper-based feature selection algorithm designed to identify all truly relevant variables in a dataset. By comparing the importance of actual features with that of randomly generated “shadow” features, Boruta systematically filters out irrelevant variables while preserving essential predictors. This rigorous selection process is particularly valuable in high-dimensional applications, such as gene expression classification, where capturing meaningful signals is crucial. For clarity, we formally define Boruta as follows:
@@
    Iteration: Remove irrelevant and tentative features and repeat until all features are decisively classified.
@@
After applying the Boruta algorithm, we retain only the relevant features (confirmed) and excluded the tentative and irrelevant features (rejected). To further enhance the selection of features in $X^{*}$, we employed the AI explanation technique outlined in the following section.
@@ 2.3 XAI for Feature Selection
Explainable AI (XAI) [13, 38] represents a forefront of AI research, aiming to elucidate the decision-making processes of complex models. In the context of gene expression classification, where feature selection is pivotal to model performance and interpretability, our study leverages LIME—Local Interpretable Model-Agnostic Explanations—to demystify and select critical features. LIME approximates the behavior of a sophisticated, black-box model with a simpler, locally interpretable surrogate, thereby pinpointing the most influential predictors in the vicinity of a given instance. This approach enhances the transparency of the model’s predictions and facilitates a more informed and rigorous feature selection process, ultimately contributing to both improved accuracy and trustworthiness of the classification system. Now, we provide a formal definition of LIME as follows:
@@
To clarify, our choice of LIME for feature selection arises from the critical question of determining the optimal number of features for the model. In this context, assessing the local importance of each vector proves to be the most effective strategy, leading us to introduce the BOLIMES algorithm. The following section will provide a comprehensive explanation of the BOLIMES algorithm and its application.
3 BOLIMES algorithm
Our methodology begins with applying the Boruta algorithm to sift through the high-dimensional gene expression dataset, effectively filtering out irrelevant features and isolating those that are truly significant. This initial reduction is critical because LIME, our subsequent interpretability tool, involves generating numerous perturbed samples and calculating distances—a computationally intensive process, especially in large feature spaces. By narrowing the feature set, we improve LIME’s precision and enhance the model’s interpretability, ultimately boosting classification performance.
@@ 14                   $f_{\text{opt}}\leftarrow f_{k}$
Algorithm 1 presents an optimal feature selection framework by integrating Boruta and LIME to refine feature subsets for classification. Given a dataset $D=(X,y)$, it first applies Boruta to extract a subset $X^{*}$ of relevant features. LIME then computes local importance scores $\mathcal{I}$ for $X^{*}$, producing a ranked set $X^{*}_{R}$ where
@@
Note that we begin with $k=10$ and increase it until $|X^{*}|$, as using a smaller number of vectors would be inefficient. The best-performing classifier defines $k^{*}$, yielding the final subset $X_{\text{opt}}$ and trained model $f_{\text{opt}}$.
@@ 4 Experiment and results
This section offers a brief description of the gene expression datasets while delivering a detailed comparative analysis of the classification models. Furthermore, we also provide the results of Boruta. Our source code has been made publicly accessible on GitHub [^2] https://github.com/Pbchung75/BOLIMES/.
@@ 4.1 Dataset and Configurations
We carry out a series of experiments using 14 gene expression datasets, all meticulously sourced from the reputable ArrayExpress repository [4]. The gene expression datasets [10] summarized in Table 1 exemplify the inherent challenges of high-dimensional biomedical data. With sample sizes ranging from $53$ to $575$ and feature counts spanning from approximately $11,950$ to over $54,600$, these datasets present a significant imbalance between the number of available samples and the vast dimensionality of gene expression profiles. Additionally, the variability in the number of classes—from as few as $3$ to as many as $10$—further complicates the classification task by introducing diverse and complex biological signals. This high dimensionality coupled with limited sample sizes accentuates the risk of overfitting and underscores the critical need for effective feature selection. In this study, our feature selection strategy is specifically designed to address these challenges, ensuring that only the most relevant features are retained for subsequent classification tasks.
@@ 14 E-GEOD-44077 226 33252 4
To evaluate our approach and the training model, we implement the AI libraries (i.e., Pandas (version 2.2.2), Scikit-learn (version 1.6.1)) and run the experiments on the computer with the following configuration: Intel Core i5-12400, 2.50 GHz, 24 GB RAM, Windows 11 Pro OS.
@@ 4.2 Boruta’s Feature Selection
In this study, we adopt Boruta as our primary feature selection method to substantially reduce the dimensionality of our gene expression datasets for the global problem. The rationale behind this choice is to prevent an explosion in computational complexity and potential loss of interpretability when LIME is applied to an excessively high-dimensional feature space. By using Boruta, we are able to effectively eliminate irrelevant features (tentative and rejected in Table 2) while retaining those that are truly informative for classification. As evidenced in Table 2, datasets such as E-GEOD-20685 are reduced from 54,627 dimensions to only 545 confirmed features, thereby rendering the subsequent LIME analysis both feasible and efficient. To achieve this, Boruta is executed with the following parameters via the BorutaPy library: a random forest classifier (‘rf‘) with 300 estimators, a maximum of 200 iterations, an alpha value of $0.01$, a strict percentile threshold of 100, two-step feature selection enabled, and a fixed random state of 42, with verbose output enabled. These parameter settings are meticulously selected to ensure a rigorous and robust selection process, ultimately facilitating a more interpretable and high-performing classification model.
@@ 4.3 Classification results
In this study, we conduct experiments with four ML algorithms (using SVM, Random Forest (RF), XGBoost (XGB), and Gradient Boosting (GB)) to determine the method yielding the highest accuracy across various gene expression datasets (see Table 3, 4, 5, and 6). We configure an SVM using SVC with a radial basis function kernel, setting C to $100,000$, gamma to $0.001$, ensuring reproducibility with a random state of $42$. In parallel, we employe a Random Forest classifier with $200$ estimators, a maximum depth of $10$. Additionally, we implemented both an XGBClassifier and a GradientBoostingClassifier, each configured with $50$ estimators, a maximum depth of $10$, and a learning rate of $0.01$. Our comparative evaluation reveales that while the SVM excelled on datasets with fewer samples, the ensemble methods—particularly Random Forest and XGBoost—demonstrated more robust performance on high-dimensional data.
In Table 3, 4, 5, and 6, we conduct SVM, RF, and Decision Tree (DT) methodology to identify the importance of features and the number of selected features ($k$ selected dimensions) following the methodology outlined in [10] i.e., Top $k$ selected feature of E-GEOD-20685 (ID 1) is $85$, E-GEOD-20711 (ID 2) is $43$, E-GEOD-21050 (ID 3) is $275$, and others. Overall, our proposed algorithm demonstrated superior performance, achieving higher accuracy with respect to the F1-score. In certain instances, it reached $100\%$, while other solutions (SVM, RF, DT) only achieved $80-90\%$. The feature selection algorithms, such as SVM, RF, and DT, with a smaller number of features, appeared to underperform in covering the data (e.g., IDs 1, 2, 4). Our approach, however, seemed to be more effective in determining the optimal number of features needed.
@@
ID Dataset Class Method Samples Top k Classification Results Training
Features Acc Prec Rec F1 Score Time (s)
1 20685 6 ours 327 107 0.955 0.956 0.955 0.955 112.137
SVM 85 0.859 0.872 0.859 0.859 0.135
RF 85 0.890 0.902 0.890 0.889 0.147
DT 85 0.872 0.884 0.872 0.868 0.192
2 20711 5 ours 90 67 0.944 0.952 0.944 0.943 17.754
SVM 43 0.811 1.000 0.811 0.811 0.818
RF 43 0.778 1.000 0.778 0.778 0.806
DT 43 0.778 1.000 0.778 0.778 0.788
3 21050 4 ours 310 70 0.677 0.709 0.677 0.678 295.982
SVM 275 0.684 0.690 0.684 0.676 0.218
RF 275 0.681 0.681 0.681 0.666 0.227
DT 275 0.652 0.674 0.652 0.622 0.250
4 21122 7 ours 158 170 0.938 0.915 0.938 0.922 82.092
SVM 78 0.886 1.000 0.886 0.886 1.489
RF 78 0.880 1.000 0.880 0.880 1.483
DT 78 0.842 1.000 0.842 0.842 1.744
5 29354 3 ours 53 10 1.000 1.000 1.000 1.000 0.165
SVM 35 0.868 1.000 0.868 0.868 0.415
RF 35 0.830 1.000 0.830 0.830 0.422
DT 35 0.660 1.000 0.660 0.660 0.397
6 30784 3 ours 229 14 1.000 1.000 1.000 1.000 3.476
SVM 42 0.904 1.000 0.904 0.904 2.277
RF 42 0.908 1.000 0.908 0.908 2.075
DT 42 0.908 1.000 0.908 0.908 2.175
7 31312 3 ours 498 43 0.900 0.900 0.900 0.900 2.144
SVM 195 0.878 0.872 0.878 0.859 0.236
RF 195 0.878 0.864 0.878 0.857 0.225
DT 195 0.809 0.796 0.809 0.775 0.345
8 31552 3 ours 111 15 0.957 0.917 0.957 0.936 1.816
SVM 44 0.820 1.000 0.820 0.820 0.931
RF 44 0.883 1.000 0.883 0.883 0.870
DT 44 0.847 1.000 0.847 0.847 0.890
9 32537 7 ours 217 40 0.795 0.730 0.795 0.747 21.201
SVM 171 0.760 1.000 0.760 0.760 3.054
RF 171 0.779 1.000 0.779 0.779 2.910
DT 171 0.756 1.000 0.756 0.756 3.450
10 33315 10 ours 575 273 0.878 0.860 0.878 0.867 33.250
SVM 2321 0.873 0.893 0.873 0.855 4.174
RF 2321 0.870 0.893 0.870 0.851 4.667
DT 2321 0.781 0.845 0.781 0.752 5.534
11 37364 4 ours 94 17 0.947 0.965 0.947 0.947 2.083
SVM 140 0.809 1.000 0.809 0.809 0.793
RF 140 0.755 1.000 0.755 0.755 0.750
DT 140 0.713 1.000 0.713 0.713 0.798
12 39582 6 ours 566 251 0.877 0.889 0.877 0.880 25.058
SVM 441 0.859 0.873 0.859 0.859 0.466
RF 441 0.862 0.874 0.862 0.861 0.480
DT 441 0.790 0.812 0.790 0.786 0.764
13 39716 3 ours 53 12 1.000 1.000 1.000 1.000 0.454
SVM 118 0.887 1.000 0.887 0.887 0.408
RF 118 0.925 1.000 0.925 0.925 0.408
DT 118 0.774 1.000 0.774 0.774 0.401
14 44077 4 ours 226 18 1.000 1.000 1.000 1.000 5.220
SVM 23 0.991 1.000 0.991 0.991 1.803
RF 23 0.996 1.000 0.996 0.996 1.790
DT 23 0.987 1.000 0.987 0.987 2.005
@@
ID Dataset Class Method Samples Top k Classification Results Training
Features Acc Prec Rec F1 Score Time (s)
1 20685 6 Ours 327 48 0.970 0.971 0.970 0.970 16306.229
SVM 85 0.838 0.852 0.838 0.831 500.484
RF 85 0.874 0.890 0.874 0.872 4.706
DT 85 0.887 0.904 0.887 0.885 5.175
2 20711 5 Ours 90 10 0.833 0.841 0.833 0.830 5.396
SVM 43 0.767 1.000 0.767 0.767 25.742
RF 43 0.833 1.000 0.833 0.833 2593.662
DT 43 0.833 1.000 0.833 0.833 2588.751
3 21050 4 Ours 310 72 0.726 0.727 0.726 0.711 1507.624
SVM 275 0.732 0.776 0.732 0.696 6.371
RF 275 0.735 0.773 0.735 0.700 6.617
DT 275 0.739 0.793 0.739 0.701 6.441
4 21122 7 Ours 158 104 0.938 0.922 0.938 0.925 1039.209
SVM 78 0.880 1.000 0.880 0.880 50.215
RF 78 0.886 1.000 0.886 0.886 50.446
DT 78 0.873 1.000 0.873 0.873 53.083
5 29354 3 Ours 53 10 1.000 1.000 1.000 1.000 3.325
SVM 35 0.887 1.000 0.887 0.887 13.873
RF 35 0.811 1.000 0.811 0.811 12.891
DT 35 0.774 1.000 0.774 0.774 12.654
6 30784 3 Ours 229 12 1.000 1.000 1.000 1.000 43.917
SVM 42 0.913 1.000 0.913 0.913 80.499
RF 42 0.895 1.000 0.895 0.895 75.487
DT 42 0.908 1.000 0.908 0.908 76.522
7 31312 3 Ours 498 161 0.900 0.821 0.900 0.858 221.322
SVM 195 0.871 0.885 0.871 0.828 8.221
RF 195 0.872 0.885 0.872 0.828 7.968
DT 195 0.841 0.859 0.841 0.800 8.715
8 31552 3 Ours 111 12 1.000 1.000 1.000 1.000 20.318
SVM 44 0.865 1.000 0.865 0.865 31.843
RF 44 0.901 1.000 0.901 0.901 29.352
DT 44 0.865 1.000 0.865 0.865 30.185
9 32537 7 Ours 217 87 0.773 0.679 0.773 0.707 1238.228
SVM 171 0.779 1.000 0.779 0.779 113.423
RF 171 0.793 1.000 0.793 0.793 104.033
DT 171 0.765 1.000 0.765 0.765 110.804
10 33315 10 Ours 575 444 0.887 0.854 0.887 0.868 985.966
SVM 2321 0.835 0.866 0.835 0.808 31.634
RF 2321 0.833 0.868 0.833 0.800 31.258
DT 2321 0.797 0.840 0.797 0.756 31.688
11 37364 4 Ours 94 10 0.947 0.953 0.947 0.947 5.579
SVM 140 0.798 1.000 0.798 0.798 26.410
RF 140 0.809 1.000 0.809 0.809 25.745
DT 140 0.734 1.000 0.734 0.734 27.083
12 39582 6 Ours 566 189 0.877 0.890 0.877 0.876 304.137
SVM 441 0.820 0.836 0.820 0.819 13.926
RF 441 0.832 0.847 0.832 0.831 13.629
DT 441 0.807 0.831 0.807 0.805 14.261
13 39716 3 Ours 53 15 1.000 1.000 1.000 1.000 19.338
SVM 118 0.925 1.000 0.925 0.925 13.133
RF 118 0.925 1.000 0.925 0.925 13.049
DT 118 0.830 1.000 0.830 0.830 13.391
14 44077 4 Ours 226 52 1.000 1.000 1.000 1.000 642.643
SVM 23 0.991 1.000 0.991 0.991 61.789
RF 23 0.996 1.000 0.996 0.996 60.206
DT 23 0.996 1.000 0.996 0.996 63.950